Last year, Sam had 30,000 to invest. He invested some of it in an account that paid 5% simple interest per year, and he invested the rest in an account that paid 8% simple interest per year. After one year, he received a total of 2040 in interest. How much did he invest in each account?
Answer by mananth(16946)

Account I 5.00% per annum ---x
Account II 8.00% per annum ---y

x + 1 y= 30000 ------------------------1
5.00% x + 8.00% y= 2040
Multiply by 100
5 x + 8 y= 204000.00 --------2
Multiply (1) by -5
we get
-5 x -5 y= -150000.00
Add this to (2)
0 x 3 y= 54000
divide by 3
y = 18000 Account I 8.00%
Balance 12000 Account II 5.00%
CHECK
18000 --------- 8.00% ------- 1440.00
12000 ------- 5.00% ------- 600.00
Total -------------- 2040.00

m.ananth@hotmail.ca